# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ Appendix C. Explicit computations of the 4-point function
where the factor of $1/k_{23}=2/s_{23}$ comes from the integration of the Koba-Nielsen factor. The calculation of this correlator will make use of the equations of motion for $U_{1}^{(3)}$, given by (2.23) and (2.24). The first term in the correlator gives
$\eqalignno{\left\langle\partial_{a}U_{1}^{(3)}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{1}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle+3\left\langle[Q,C_{1,a}]\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle\cr&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{1}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}\{Q,\tilde{\phi}_{23}^{a}\}U_{4}^{(3)}\right\rangle.&({\hbox{C.}2})}$
@@
$\eqalignno{{\cal A}_{4}^{u}&={2\over s_{23}}\Bigg{(}3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{1}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}(\{Q,\tilde{\phi}_{23}^{a}\}-(\lambda\gamma^{a}\tilde{\phi}_{23}))U_{4}^{(3)}\right\rangle\cr&\quad\quad\quad-3\left\langle C_{1,\alpha}([Q,\tilde{\phi}_{23}^{\alpha}]-\tilde{\Omega}^{\alpha}_{23})U_{4}^{(3)}\right\rangle\Bigg{)}.&({\hbox{C.}5})}$
@@
where we contract $U^{(3)}$ with $U^{(0)}$ using eqn. (2.45), with $C_{ij}=C_{i,a}\phi_{j}^{a}+C_{i,\alpha}\phi_{j}^{\alpha}$ and $s_{ij}=(k_{i}+k_{j})^{2}=2k_{ij}$. Now, let us apply the equations of motion for $U_{12}^{(3)}$ (2.48) and (2.49). The first term of eqn. (C.6) becomes
$\eqalignno{\left\langle\partial_{a}U_{12}^{(3)}\phi_{3}^{a}U_{4}^{(3)}\right\rangle&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle[Q,C_{12,a}]\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3k_{12}\left\langle M_{12,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle C_{12,a}\{Q,\phi_{3}^{a}\}U_{4}^{(3)}\right\rangle+3k_{12}\left\langle M_{12,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle,\cr&&({\hbox{C.}7})}$
where $M_{12,a}=(\lambda\gamma_{ab}\phi_{(1})\phi_{2)}^{b}+(\lambda\gamma_{bc}\lambda)\phi_{(1}^{c}h^{b}_{2),a}$. The second term reads
$\eqalignno{\left\langle D_{\alpha}U_{12}^{(3)}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle&=-3\left\langle\{Q,C_{12,\alpha}\}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma^{a})_{\alpha}C_{12,a}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+3k_{12}\left\langle M_{12,\alpha}^{(3)}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle\cr&=-3\left\langle C_{12,\alpha}[Q,\phi_{3}^{\alpha}]U_{4}^{(3)}\right\rangle-3\left\langle C_{12,a}(\lambda\gamma^{a}\phi_{3})U_{4}^{(3)}\right\rangle+3k_{12}\left\langle M_{12,\alpha}^{(3)}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle,\cr&&({\hbox{C.}8})}$
where $M_{12,\alpha}=2(\lambda\gamma_{ab})_{\alpha}\phi^{a}_{1}\phi_{2}^{b}+(\lambda\gamma_{ab}\lambda)h_{(1,\alpha}^{a}\phi_{2)}^{b}$. Finally, the third term gives
@@
$\eqalignno{{\cal A}_{4}^{s}&={6\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{b}_{2}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{1}^{c}h^{b}_{2,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{bc}\lambda)\phi_{2}^{c}h^{b}_{1,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{2}^{a}h_{1,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{2,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle.&({\hbox{C.}10})}$
@@
$\eqalignno{&3\left\langle C_{1a}\{\phi^{a}_{2},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+3\left\langle C_{1\alpha}[\phi^{\alpha}_{2},U_{3}^{(1)}]U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr={}&3\left\langle C_{1a}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big{)}U_{4}^{(3)}\right\rangle\cr&-3\left\langle C_{1\alpha}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big{)}U_{4}^{(3)}\right\rangle.&({\hbox{C.}12})}$
@@
$\eqalignno{&3\left\langle\phi^{a}_{2}[C_{1a},U_{3}^{(1)}]U_{4}^{(3)}\right\rangle-3\left\langle\phi^{\alpha}_{2}\{C_{1\alpha},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr=&3\left\langle(\phi^{a}_{(2}\phi_{3)}^{c}\partial_{c}C_{1a}+\phi^{a}_{(2}\phi_{3)}^{\alpha}D_{\alpha}C_{1a}+\phi^{a}_{(2}\Omega_{3)}^{\alpha}C_{1,\alpha a})U_{4}^{(3)}\right\rangle\cr=&6\left\langle(\lambda\gamma_{ab}\phi_{1})\phi_{2}^{a}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{2}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\cr&-3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{c}h_{1,c}^{b}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{\alpha}h_{1,\alpha}^{b}U_{4}^{(3)}\right\rangle\cr&+3\left\langle\left\{Q,\Big{(}\phi^{a}_{(2}\phi^{b}_{3)}C_{1,ab}+\phi^{a}_{(2}\phi^{\alpha}_{3)}C_{1,a\alpha}+\phi^{\alpha}_{(2}\phi^{\beta}_{3)}C_{1,\alpha\beta}\Big{)}U_{4}^{(3)}\right\}\right\rangle.&({\hbox{C.}13})}$
@@
$\eqalignno{&3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr=&6\left\langle(\lambda\gamma_{ab}\phi_{1})\phi_{2}^{a}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{2}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\cr&-3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{c}h_{1,c}^{b}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{\alpha}h_{1,\alpha}^{b}U_{4}^{(3)}\right\rangle\cr&+3\left\langle C_{1a}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big{)}U_{4}^{(3)}\right\rangle\cr&-3\left\langle C_{1\alpha}\Big{(}\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big{)}U_{4}^{(3)}\right\rangle\cr&+3\left\langle\left\{Q,\Big{(}\phi^{a}_{(2}\phi^{b}_{3)}C_{1,ab}+\phi^{a}_{(2}\phi^{\alpha}_{3)}C_{1,a\alpha}+\phi^{\alpha}_{(2}\phi^{\beta}_{3)}C_{1,\alpha\beta}\Big{)}U_{4}^{(3)}\right\}\right\rangle.&({\hbox{C.}14})}$
Combining this with the expressions (3.16) and (3.18), one obtains the result given in (3.19).
@@ C.1. Compact form for the 4-point function
$\eqalignno{&\partial_{a}U^{(3)}-3[Q,C_{a}]=3(\lambda\gamma_{ab}\lambda)\phi^{b},&({\hbox{C.}15})\cr&D_{\alpha}U^{(3)}+3\{Q,C_{\alpha}\}=-3(\lambda\gamma^{a})_{\alpha}C_{a},&({\hbox{C.}16})\cr&\partial_{a}U^{(3)}_{12}-3[Q,C_{a,12}]=3(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}+3k_{12}M_{a,12},&({\hbox{C.}17})\cr&D_{\alpha}U_{12}^{(3)}+3\{Q,C_{\alpha,12}\}=-3(\lambda\gamma^{a})_{\alpha}C_{a,12}+3k_{12}M_{\alpha,12},&({\hbox{C.}18})}$
@@
$\eqalignno{&\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=-{1\over 3}\left\langle\partial_{a}U^{(3)}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+\left\langle[Q,C_{a,12}]\phi_{3}^{a}U_{4}^{(3)}\right\rangle+k_{12}\left\langle M_{a,12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr={}&{1\over 3}\left\langle U^{(3)}_{12}\phi_{3}^{a}\partial_{a}U_{4}^{(3)}\right\rangle-\left\langle C_{a,12}\{Q,\phi_{3}^{a}\}U_{4}^{(3)}\right\rangle+k_{12}\left\langle M_{a,12}\phi^{a}_{3}U_{4}^{(3)}\right\rangle\cr={}&\left\langle U^{(3)}_{12}(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}\phi_{4}^{b}\right\rangle+\left\langle U_{12}^{(3)}\phi_{3}^{a}[Q,C_{a,4}]\right\rangle-\left\langle C_{a,12}(\lambda\gamma^{a}\phi_{3}))U_{4}^{(3)}\right\rangle+k_{12}\left\langle M_{a,12}\phi^{a}_{3}U_{4}^{(3)}\right\rangle,\cr&&({\hbox{C.}20})}$
@@
$\eqalignno{&\left\langle U^{(3)}_{12}(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}\phi_{4}^{b}\right\rangle={2\over 3}\left\langle U_{12}^{(3)}\left\{Q,\left(U_{34}^{(3)}/s_{34}\right)\right\}\right\rangle={2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{12}^{(3)}\}\right\rangle.&({\hbox{C.}21})}$
Having identified this important peace of the computation, next we integrate $Q$ by parts in the second term of the last line in (C.20). We get
$\eqalignno{\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=&{2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{34}^{(3)}\}\right\rangle+\left\langle U_{12}^{(3)}(\lambda\gamma^{a}\phi_{3})C_{a,4}\right\rangle-\left\langle C_{a,12}(\lambda\gamma^{a}\phi_{3})U_{4}^{(3)}\right\rangle\cr&{}-3k_{12}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\phi_{3}^{e}C_{e,4}\right\rangle+k_{12}\left\langle M_{a,12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle.&({\hbox{C.}22})}$
@@
$\eqalignno{\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=&{2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{34}^{(3)}\}\right\rangle+k_{12}\left\langle(M_{a,12}\phi_{3}^{a}-M_{\alpha,12}\phi_{3}^{\alpha})U_{4}^{(3)}\right\rangle\cr&-3k_{12}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}(\phi_{3}^{e}C_{e,4}+\phi_{3}^{\alpha}C_{\alpha,4})\right\rangle.&({\hbox{C.}23})}$
We will now write $C_{a}$ and $C_{\alpha}$ in terms of the physical operators ${\tenbf C}_{a}$ and ${\tenbf C}_{\alpha}$ defined in (2.14) and (2.13)

### v2

@@ Appendix C. Explicit computations of the 4-point function
where the overall factor $2/s_{23}$ comes from the integration of the Koba-Nielsen measure. The calculation of this correlator will make use of the equations of motion for $U_{1}^{(3)}$, given by (2.23) and (2.24). The first term in the correlator gives
$\eqalignno{\left\langle\partial_{a}U_{1}^{(3)}\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\tilde{\phi}_{23}^{b}U_{4}^{(3)}\right\rangle+3\left\langle[Q,C_{1,a}]\tilde{\phi}_{23}^{a}U_{4}^{(3)}\right\rangle\cr&=3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\tilde{\phi}_{23}^{b}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}\{Q,\tilde{\phi}_{23}^{a}\}U_{4}^{(3)}\right\rangle.&({\hbox{C.}2})}$
@@
$\eqalignno{{\cal A}_{4}^{u}&={2\over s_{23}}\Bigg(3\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\tilde{\phi}_{23}^{b}U_{4}^{(3)}\right\rangle+3\left\langle C_{1,a}(\{Q,\tilde{\phi}_{23}^{a}\}-(\lambda\gamma^{a}\tilde{\phi}_{23}))U_{4}^{(3)}\right\rangle\cr&\quad\quad\quad-3\left\langle C_{1,\alpha}([Q,\tilde{\phi}_{23}^{\alpha}]-\tilde{\Omega}^{\alpha}_{23})U_{4}^{(3)}\right\rangle\Bigg).&({\hbox{C.}5})}$
@@
where we contract $U^{(3)}$ with $U^{(0)}$ using eqn. (2.45), with $C_{ij}=C_{i,a}\phi_{j}^{a}+C_{i,\alpha}\phi_{j}^{\alpha}$. Now, let us apply the equations of motion for $U_{12}^{(3)}$ (2.48) and (2.49). The first term of eqn. (C.6) becomes
$\eqalignno{\left\langle\partial_{a}U_{12}^{(3)}\phi_{3}^{a}U_{4}^{(3)}\right\rangle&=3\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle[Q,C_{12,a}]\phi_{3}^{a}U_{4}^{(3)}\right\rangle+{3s_{12}\over 2}\left\langle M_{12,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr&=3\left\langle(\lambda\gamma_{ba}\lambda)\phi^{b}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle C_{12,a}\{Q,\phi_{3}^{a}\}U_{4}^{(3)}\right\rangle+{3s_{12}\over 2}\left\langle M_{12,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle,\cr&\quad&({\hbox{C.}7})}$
where $M_{12,a}=2(\lambda\gamma_{ab}\phi_{(1})\phi_{2)}^{b}+2(\lambda\gamma_{bc}\lambda)\phi_{(1}^{c}h^{b}_{2),a}$. The second term reads
$\eqalignno{\left\langle D_{\alpha}U_{12}^{(3)}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle&=-3\left\langle\{Q,C_{12,\alpha}\}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle-3\left\langle(\lambda\gamma^{a})_{\alpha}C_{12,a}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+{3s_{12}\over 2}\left\langle M_{12,\alpha}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle\cr&=-3\left\langle C_{12,\alpha}[Q,\phi_{3}^{\alpha}]U_{4}^{(3)}\right\rangle-3\left\langle C_{12,a}(\lambda\gamma^{a}\phi_{3})U_{4}^{(3)}\right\rangle+{3s_{12}\over 2}\left\langle M_{12,\alpha}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle,\cr&\quad&({\hbox{C.}8})}$
where $M_{12,\alpha}=2(\lambda\gamma_{ab})_{\alpha}\phi^{a}_{1}\phi_{2}^{b}+2(\lambda\gamma_{ab}\lambda)h_{(1,\alpha}^{a}\phi_{2)}^{b}$. Finally, the third term gives
@@
$\eqalignno{{\cal A}_{4}^{s}&={6\over s_{12}}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{12}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{ba}\phi_{1})\phi^{b}_{2}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ba}\phi_{2})\phi^{b}_{1}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+6\left\langle(\lambda\gamma_{ba}\phi_{3})\phi^{b}_{1}\phi_{2}^{a}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{1}^{c}h^{b}_{2,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{cb}\lambda)\phi_{2}^{c}h^{b}_{1,a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr&+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{2}^{a}h_{1,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\lambda)\phi_{1}^{a}h_{2,\alpha}^{b}\phi_{3}^{\alpha}U_{4}^{(3)}\right\rangle.&({\hbox{C.}10})}$
@@
$\eqalignno{&3\left\langle C_{1a}\{\phi^{a}_{2},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+3\left\langle C_{1\alpha}[\phi^{\alpha}_{2},U_{3}^{(1)}]U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr={}&6\left\langle C_{1a}\Big(\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big)U_{4}^{(3)}\right\rangle\cr&-6\left\langle C_{1\alpha}\Big(\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big)U_{4}^{(3)}\right\rangle.&({\hbox{C.}12})}$
@@
$\eqalignno{&3\left\langle\phi^{a}_{2}[C_{1a},U_{3}^{(1)}]U_{4}^{(3)}\right\rangle-3\left\langle\phi^{\alpha}_{2}\{C_{1\alpha},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr=&6\left\langle(\phi^{a}_{(2}\phi_{3)}^{c}\partial_{c}C_{1a}+\phi^{a}_{(2}\phi_{3)}^{\alpha}D_{\alpha}C_{1a}+\phi^{a}_{(2}\Omega_{3)}^{\alpha}C_{1,\alpha a})U_{4}^{(3)}\right\rangle\cr=&6\left\langle(\lambda\gamma_{ab}\phi_{1})\phi_{2}^{a}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{2}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\cr&-6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{c}h_{1,c}^{b}U_{4}^{(3)}\right\rangle-6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{\alpha}h_{1,\alpha}^{b}U_{4}^{(3)}\right\rangle\cr&+6\left\langle\left\{Q,\Big(\phi^{a}_{(2}\phi^{b}_{3)}C_{1,ab}+\phi^{a}_{(2}\phi^{\alpha}_{3)}C_{1,a\alpha}+\phi^{\alpha}_{(2}\phi^{\beta}_{3)}C_{1,\alpha\beta}\Big)U_{4}^{(3)}\right\}\right\rangle.&({\hbox{C.}13})}$
@@
$\eqalignno{&3\left\langle\{C_{12},U_{3}^{(1)}\}U_{4}^{(3)}\right\rangle+(2\leftrightarrow 3)\cr=&6\left\langle(\lambda\gamma_{ab}\phi_{1})\phi_{2}^{a}\phi_{3}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{2})\phi_{3}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle+3\left\langle(\lambda\gamma_{ab}\phi_{3})\phi_{2}^{a}\phi_{1}^{b}U_{4}^{(3)}\right\rangle\cr&-6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{c}h_{1,c}^{b}U_{4}^{(3)}\right\rangle-6\left\langle(\lambda\gamma_{ab}\lambda)\phi_{(2}^{a}\phi_{3)}^{\alpha}h_{1,\alpha}^{b}U_{4}^{(3)}\right\rangle\cr&+6\left\langle C_{1a}\Big(\phi_{(2}^{c}\partial_{c}\phi^{a}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{a}_{3)}-\Omega_{(2}^{\alpha}h^{a}_{3),\alpha}\Big)U_{4}^{(3)}\right\rangle\cr&-6\left\langle C_{1\alpha}\Big(\phi_{(2}^{c}\partial_{c}\phi^{\alpha}_{3)}+\phi_{(2}^{\delta}D_{\delta}\phi^{\alpha}_{3)}-\Omega_{(2}^{\beta}h^{\alpha}_{3),\beta}\Big)U_{4}^{(3)}\right\rangle\cr&+6\left\langle\left\{Q,\Big(\phi^{a}_{(2}\phi^{b}_{3)}C_{1,ab}+\phi^{a}_{(2}\phi^{\alpha}_{3)}C_{1,a\alpha}+\phi^{\alpha}_{(2}\phi^{\beta}_{3)}C_{1,\alpha\beta}\Big)U_{4}^{(3)}\right\}\right\rangle.&({\hbox{C.}14})}$
Combining this with the expressions (3.17) and (3.19), one obtains the result given in eqn. (3.20).
@@ C.1. Compact form for the 4-point function
$\eqalignno{&\partial_{a}U^{(3)}-3[Q,C_{a}]=3(\lambda\gamma_{ab}\lambda)\phi^{b},&({\hbox{C.}15})\cr&D_{\alpha}U^{(3)}+3\{Q,C_{\alpha}\}=-3(\lambda\gamma^{a})_{\alpha}C_{a},&({\hbox{C.}16})\cr&\partial_{a}U^{(3)}_{12}-3[Q,C_{a,12}]=3(\lambda\gamma_{ab}\lambda)\phi^{b}_{12}+{3s_{12}\over 2}M_{a,12},&({\hbox{C.}17})\cr&D_{\alpha}U_{12}^{(3)}+3\{Q,C_{\alpha,12}\}=-3(\lambda\gamma^{a})_{\alpha}C_{a,12}+{3s_{12}\over 2}M_{\alpha,12},&({\hbox{C.}18})}$
@@
$\eqalignno{&\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=-{1\over 3}\left\langle\partial_{a}U^{(3)}_{12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle+\left\langle[Q,C_{a,12}]\phi_{3}^{a}U_{4}^{(3)}\right\rangle+{s_{12}\over 2}\left\langle M_{a,12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle\cr={}&{1\over 3}\left\langle U^{(3)}_{12}\phi_{3}^{a}\partial_{a}U_{4}^{(3)}\right\rangle-\left\langle C_{a,12}\{Q,\phi_{3}^{a}\}U_{4}^{(3)}\right\rangle+{s_{12}\over 2}\left\langle M_{a,12}\phi^{a}_{3}U_{4}^{(3)}\right\rangle\cr={}&\left\langle U^{(3)}_{12}(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}\phi_{4}^{b}\right\rangle+\left\langle U_{12}^{(3)}\phi_{3}^{a}[Q,C_{a,4}]\right\rangle-\left\langle C_{a,12}(\lambda\gamma^{a}\phi_{3}))U_{4}^{(3)}\right\rangle+{s_{12}\over 2}\left\langle M_{a,12}\phi^{a}_{3}U_{4}^{(3)}\right\rangle,\cr&&({\hbox{C.}20})}$
@@
$\eqalignno{&\left\langle U^{(3)}_{12}(\lambda\gamma_{ab}\lambda)\phi_{3}^{a}\phi_{4}^{b}\right\rangle={2\over 3}\left\langle U_{12}^{(3)}\left\{Q,\left(U_{34}^{(3)}/s_{34}\right)\right\}\right\rangle={2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{34}^{(3)}\}\right\rangle.&({\hbox{C.}21})}$
Having identified this important piece of the computation, next we integrate $Q$ by parts in the second term of the last line in (C.20). We get
$\eqalignno{\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=&{2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{34}^{(3)}\}\right\rangle+\left\langle U_{12}^{(3)}(\lambda\gamma^{a}\phi_{3})C_{a,4}\right\rangle-\left\langle C_{a,12}(\lambda\gamma^{a}\phi_{3})U_{4}^{(3)}\right\rangle\cr&{}-{3s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\phi_{3}^{e}C_{e,4}\right\rangle+{s_{12}\over 2}\left\langle M_{a,12}\phi_{3}^{a}U_{4}^{(3)}\right\rangle.&({\hbox{C.}22})}$
@@
$\eqalignno{\left\langle(\lambda\gamma_{ab}\lambda)\phi_{12}^{a}\phi_{3}^{a}U_{4}^{(3)}\right\rangle=&{2\over 3}\left\langle U_{12}^{(3)}\{Q,V_{34}^{(3)}\}\right\rangle+{s_{12}\over 2}\left\langle(M_{a,12}\phi_{3}^{a}-M_{\alpha,12}\phi_{3}^{\alpha})U_{4}^{(3)}\right\rangle\cr&-{3s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}(\phi_{3}^{e}C_{e,4}+\phi_{3}^{\alpha}C_{\alpha,4})\right\rangle.&({\hbox{C.}23})}$
We will now write $C_{a}$ and $C_{\alpha}$ in terms of the physical operators ${\tenbf C}_{a}$ and ${\tenbf C}_{\alpha}$ defined in eqns. (2.13)-(2.14)